### v1

# Foundation Models Meet Low-Cost Sensors: Test-Time Adaptation for Rescaling Disparity for Zero-Shot Metric Depth Estimation

Abstract
The recent development of foundation models for monocular depth estimation such as Depth Anything paved the way to zero-shot monocular depth estimation. Since it returns an affine-invariant disparity map, the favored technique to recover the metric depth consists in fine-tuning the model. However, this stage is costly to perform because of the training but also due to the creation of the dataset. It must contain images captured by the camera that will be used at test time and the corresponding ground truth. Moreover, the fine-tuning may also degrade the generalizing capacity of the original model. Instead, we propose in this paper a new method to rescale Depth Anything predictions using 3D points provided by low-cost sensors or techniques such as low-resolution LiDAR, stereo camera, structure-from-motion where poses are given by an IMU. Thus, this approach avoids fine-tuning and preserves the generalizing power of the original depth estimation model while being robust to the noise of the sensor or of the depth model. Our experiments highlight improvements relative to other metric depth estimation methods and competitive results compared to fine-tuned approaches. Code available at gitlab.ensta.fr/ssh/monocular-depth-rescaling.
I Introduction
Despite the growth of 3D perception sensors such as LiDAR, time-of-flight, structured light or stereo cameras, for robotic systems, traditional monocular cameras remain a key sensor for any robot setup. In addition to being a cheaper solution, monocular depth estimation can also offer denser outputs as well as larger depth range. The increase in the number of open datasets for monocular depth estimation and the development of neural network architectures such as Vision Transformers that scale well with the size of the training dataset [1] allowed the emergence of foundation models for monocular depth estimation [2].
However, predicting the depth from a single image is an ambiguous task since depth depends on the image content and on the camera calibration. The consequence is that models trained on a single dataset to make metric depth (also referred to as absolute depth) predictions generalize poorly on images captured with a different camera. To avoid this issue during training on multiple datasets, some methods [3, 2, 4] learn an affine-invariant depth or disparity which allows impressive results on zero-shot relative depth estimation benchmarks. The most common way of adapting these models in order to make metric depth predictions consists in fine-tuning on the target domain, i.e., on a dataset composed of images captured with the target camera calibration. Such a solution is costly due to the creation of the dataset and the training computation. Moreover, transferring performance on images with a different calibration still remains impossible. Several solutions have been proposed to solve this issue by taking into account the camera calibration in the method. However, such approaches are often more costly at inference or cannot be trained on image datasets with unknown calibration such as ImageNet [5] contrary to methods like Depth Anything [2, 4].
In this paper, we investigate the use of low-cost sensors for solving the scale ambiguity in order to perform zero-shot metric depth estimation given affine-invariant disparity predictions. The additional sensor provides some reference 3D points that are exploited to recover the scaling parameters. We note for the sake of brevity, we will refer to rescaling the process of finding an affine transformation and applying it to recover the metric depth. We explored several low-cost sensors including low-resolution LiDARs (with 16 and 32 beams) as well as 2D LiDARs (with a single beam) that are often used for indoor robotics, stereo cameras and IMU. The latter provides the relative camera pose that is leveraged to triangulate matching points in successive images in a structure-from-motion strategy. The advantage of such an approach is twofold. On the one hand, it can be used with any monocular depth estimation model such as Depth Anything V1 and V2 [2, 4] with a high generalization ability due to their large and diverse training dataset. On the other hand, our method does not require any costly fine-tuning on the target domain and provides instant adaptation. We conducted extensive experiments that demonstrate the interest of our approaches on standard metric depth estimation benchmarks.
II Related work
Figure 1: Illustration of our method. First, an affine-invariant disparity map $d$ is predicted from an image with a neural network such as Depth Anything [2] while in parallel a low-cost sensor is used to estimate a set of 3D points $P$. $P$ and the corresponding values in $d$ are then used to estimate the scaling parameters $\alpha$ and $\beta$ using a RANSAC [6]. The parameters are applied to $d$ to recover the metric depth $D$.
II-A Monocular depth estimation
While pioneer works on monocular depth estimation relied on Markov Random Fields [7, 8], subsequent ones have shown the effectiveness of convolutional neural networks [9, 10] then transformers for this task [11]. More recently, benefiting from advances in image generation [12], impressive results have been obtained using diffusion models [13, 14]. Regarding the output, monocular depth estimation has been initially addressed as a regression problem [9], before moving to classification approaches with discrete bins that showed better performances [15, 16, 17]. Monocular depth estimation methods can also be divided whether they are supervised [9, 10, 15] or self-supervised [18]. In the latter case, depth is learned with a photometric loss that compares a target image with its reconstruction obtained from another view of the same scene, the depth prediction and the camera pose. Recently, the multiplication of depth estimation benchmark allowed training models on multiple domains at once [3] and paved the way to zero-shot monocular depth estimation. Due to the inherent scale ambiguity of depth estimation from a single image, these methods are mostly trained to produce affine-invariant depth or disparity predictions [3, 11, 2]. Our work aims to be used with any depth estimation model as long as it returns disparity maps that are accurate within an affine transformation whatever its other characteristics in terms of training set or architecture.
II-B Scale estimation for monocular depth
A major issue of monocular depth prediction relies on scale ambiguity which means that the true size of an object cannot be recovered for sure from a single image. Since most monocular depth estimation models are trained for a specific camera calibration, using them with another camera leads to ill-scaled predictions. Likewise, self-supervised depth estimation methods may also predict depth up to an affine transformation if they are trained with no reference of absolute distance at all (most of the time when the camera pose is also estimated by a neural network). The most common way to recover the metric depth in this situation is to fine-tune the model on a dataset collected with the camera that will be used at inference. However, this solution is costly to implement because it needs the creation of an image dataset with the relative ground truth and a new training. In case of self-supervised depth estimation, the scaling factor can be recovered from the ground truth pose given by an IMU to compute the photometric loss [19], the constraints of a stereo camera [20, 18, 21] or the speed of the vehicle that carries the camera [22, 23]. Other works focus on the temporal consistency of the scale of depth predictions [24, 25]. Closer to our work, [26] learns to predict scaled depth maps from affine-invariant disparity maps and a visual-inertial odometry. In [27], authors propose to estimate the scale factor for a target domain from a model that has been jointly trained in a self-supervised manner on a source domain with known ground truth and on a target domain without depth annotation. In contrast, we propose to rescale any disparity prediction that is correct up to an affine transformation with no training or fine-tuning at all but by leveraging reference 3D points provided by an external low-cost sensor. Furthermore, the use of an external sensor makes our approach adaptable to any camera calibration.
II-C Monocular zero-shot metric depth estimation
ZoeDepth [28] is the first zero-shot metric depth estimation method. It first consists in relative depth pre-training of the MiDaS [3] backbone then fine-tuning two metric bins modules, one for indoor scenes and the other for outdoor scenes. More recently, several zero-shot metric depth estimation methods have been proposed without any fine-tuning. Thus, ScaleDepth [29] decomposes metric depth estimation in relative depth estimation and scale estimation each with a dedicated module. It can also leverage a text description of the scene to guide the supervision. In [30], authors introduce a dedicated architecture that takes as input the calibration matrix in addition to the image. On the contrary, [31] makes predictions from the images only but applies transformation on the input images so the predictions are invariant to the image size or the camera calibration. UniDepth [32] estimates an internal representation of the camera calibration directly from the input images only. These approaches still have drawbacks since they are often more costly at inference or need for the calibration of the images even at the training stage. This prevents exploiting image datasets with unknown calibration as Depth Anything V1 [2] and V2 [4] which leverage a distillation strategy with such datasets to improve their generalization abilities.
III Method
Let $\phi$ be a monocular depth estimation model such as [3, 2] that is trained to predict an affine-invariant disparity map $d\in\mathbb{R}^{H\times W}$ given an input image $I\in\mathbb{R}^{H\times W\times 3}$ where $H$ and $W$ are the height and the width of the image $I$. Therefore, the metric or absolute depth map $D\in\mathbb{R}^{H\times W}$ that corresponds to the inverse of the metric or absolute disparity $D^{-1}$ given by the relation:
$D^{-1}=\alpha d+\beta,$ (1)
where parameters $\alpha\in\mathbb{R}^{+}_{*}$ and $\beta\in\mathbb{R}$ are the unknown scaling factor and the offset, respectively.
Our method, illustrated in Fig. 1, aims at recovering the metric depth map $D$ from the affine-invariant disparity map $d$ and a set of $N$ reference 3D points $P\in\mathbb{R}^{N\times 3}$ by regressing the parameters $\alpha$ and $\beta$. First, we perform a bilinear sampling on the affine-invariant map $d$ at the locations of the projection of the reference 3D points $P$ on the image plan so as to have an affine-invariant disparity value for each reference 3D point. Second, we leverage a linear regression to estimate parameters $\alpha$ and $\beta$. To favor the robustness to potential outliers in the reference 3D points, we use a RANSAC algorithm [6], but it could be replaced by another robust regression method.
In our approach, we assume the set $P$ is provided with a low-cost sensor like a low-resolution LiDAR, a stereo camera or an IMU from which the poses of the camera that capture $I$ can be extracted. The poses are used in a structure-from-motion (SFM) strategy to triangulate matching points in consecutive images of a video. Thus, $P$ may be a very sparse depth map, i.e., $N\ll HW$ especially when provided by a low-resolution LiDAR or by a structure-from-motion technique. It may also contain some outlier measures for instance the depth of the most distant pixel given by a stereo camera or when applying structure-from-motion with dynamic objects.
As our method relies on affine-invariant disparity maps, we can exploit, without any fine-tuning, foundation models like Depth Anything V1 or V2 [2, 4] that are trained to predict such outputs. Furthermore, the robustness to outliers enabled by the RANSAC allows low-quality sensors to be used to obtain the reference 3D points. Thus, we can provide low-cost metric depth maps whatever the environment or the camera. Nevertheless, since Depth Anything processes each image independently and is supervised in such a way there is no guarantee its disparity predictions are normalized identically even for consecutive images of a video, the scaling factor $\alpha$ and the offset $\beta$ must be estimated for each image.
IV Experiments
IV-A Dataset and metrics
To evaluate the performance in zero-shot metric depth estimation, we focus on standard benchmarks for monocular depth estimation that have not been used to pretrain Depth Anything [2]. This includes indoor datasets: NYU V2 [33], SUN-RGBD [34], IBIMS-1 [35] and DIODE indoor [36] and outdoor datasets: KITTI [37] , DDAD [23] and DIODE outdoor [36]. We also use DIML dataset [38] that appears in the training set of Depth Anything to show the limitations of rescaling with stereo depth. We adopt standard depth estimation metrics (see [9]) to compare our method to other approaches of the literature.
IV-B Implementation details
All experiments have been conducted using Depth Anything V1 [2] with ViT Large [1] without fine-tuning unless otherwise mentioned. We adopt Depth Anything V1 [2] code base and settings for evaluation.
We simulate low-resolution LiDARs by evenly selecting as many horizontal lines in the ground truth depth maps as the number of beams in 32-laser, 16-laser or 2D LiDARs (i.e., with a single beam). We study rescaling using stereo depth on KITTI dataset only since it provides stereo images. For this purpose, we estimate stereo depth maps using Semi Global Matching (SGM) [39] with the settings from [40]. We study the structure from motion for KITTI [37] and DDAD datasets [23] as they both consist of temporal sequences of images including the pose estimations. To obtain the reference 3D points, we first extract matching keypoints in the target image and the previous one using SIFT [41] or OmniGlue [42] and triangulate them using the pose between these images. Since this strategy requires enough displacement magnitude between the two images, we only consider image couples with a rotation higher than $5$ degrees and a translation greater than $1.5$ and $2$ meters for KITTI and DDAD datasets, respectively. We note that this threshold is the only hyperparameter that needs to be tunned in our approach.
IV-C Results
TABLE I: Quantitative results on the NYU V2 benchmark [33] (indoor).
Methods $\delta_{1}$uparrow $\delta_{2}$uparrow $\delta_{3}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ log10 $\downarrow$
ZeroDepth [30] (ZS) 0.926 0.986 0.081 0.338
Metric3D (ZS) 0.944 0.986 0.995 0.083 0.310 0.035
Metric3D V2 [31] (ZS) 0.975 0.994 0.998 0.063 0.251 0.028
[…]
ZeroDepth [30] (FT) 0.954 0.995 1.000 0.074 0.269 0.103
[…]
Ours rescaling w/ LiDAR 1 beam 0.939 0.981 0.993 0.063 0.652 0.028
[…]
TABLE II: Quantitative results on the KITTI benchmark [37] (outdoor).
Methods $\delta_{1}$uparrow $\delta_{2}$uparrow $\delta_{3}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ RMSE Log $\downarrow$
Semi-Global Matching [39] (stereo) 0.855 0.947 0.976 0.160 7.351 0.227
ZeroDepth [30] (ZS) 0.910 0.980 0.996 0.102 4.044 0.172
Metric3D (ZS) 0.964 0.993 0.998 0.058 2.770 0.092
[…]
ScaleDepth [29] (FT) 0,980 0,980 1.000 0,048 1,987 0,073
[…]
Ours rescaling w/ LiDAR 1 beam 0.891 0.974 0.986 0.131 3.096 0.142
[…]
Ours rescaling w/ SFM (SIFT [41]) 0.893 0.985 0.998 0.103 3.920 0.134
[…]
Ours rescaling w/ stereo (SGM [39]) 0.941 0.989 0.997 0.072 3.526 0.111
TABLE III: Quantitative results on different zero-shot indoor benchmarks.
Methods SUN-RGBD IBIMS-1 DIODE Indoor
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
Metric3D 0.144 0.252
Metric3D V2 [31] 0.185 0.592 0.093 0.389
Unidepth [32] 0.966 0.797 0.774
[…]
Depth Anything [2] (FT NYU V2) 0.658 0.500 0.616 0.714 0.150 0.593 0.303 0.325 1.476
[…]
TABLE IV: Quantitative results on different zero-shot outdoor benchmarks.
Methods DIODE Outdoor DDAD
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
ZoeDepth [28] (FT KITTI) 0.814 0.156 10.678
ZeroDepth [30] 0.814 0.156 10.678
[…]
Metric3D V2 [31] 0.221 3.897
Unidepth [32] 0.864
ScaleDepth [29] (FT KITTI) 0.333 0.605 6.950 0.863 0.120 6.378
Depth Anything [2] (FT KITTI) 0.288 0.794 6.641 0.886 0.105 5.931
Ours rescaling w/ Lidar 1 beam 0.689 0.880 6.222 0.706 0.326 9.229
[…]
Ours rescaling w/ SFM (SIFT [41]) 0.794 0.188 6.981
Ours rescaling w/ SFM (OmniGlue [42]) 0.947 0.112 5.300
TABLE V: Comparison of the metric depth estimated by SGM [39] or by rescaling Depth Anything predictions using SGM.
Methods KITTI DIML Outdoor
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
SGM [39] 0.855 0.16 7.351 0.973 0.118 1.503
Ours rescaling w/ stereo (SGM [39]) 0.941 0.072 3.526 0.96 0.123 1.509
TABLE VI: Cost comparison for different architectures.
Methods #param (M) runtime (ms)
Metric3D V2 [31] (ZS) 412 194
Unidepth [32] (ZS) 347 140
ZoeDepth [28] (FT) 335 113
Depth Anything [2] (FT) 335 113
Depth Anything V2[4] (FT) 336 128
Ours Depth Anything [2] + rescaling 335 120
In Tab. I, we compare our LiDAR-based rescaling approach with zero-shot metric depth estimation methods and other depth estimation methods that have been fine-tuned on the NYU V2 dataset [33]. We conduct a similar study in Tab. II on the KITTI dataset [37] with in addition results of rescaling with 3D reference points provided by structure-from-motion or by stereo depth estimation. For a fair comparison, we also include the performance of stereo depth estimation using Semi Global Matching [39] only. Tab. III and Tab. IV provide additional comparisons on indoor and outdoor datasets, respectively. For each dataset, our results show rescaling using LiDARs with different numbers of beams and for the DDAD dataset [23], we also leverage structure-from-motion. Finally, we analyze the cost of some of the methods we compare in Tab. VI. The runtimes have been measured on the same NVIDIA GeForce RTX $3090$ GPU for each approach.
Rescaling with LiDAR
Our experiments highlight an overall benefit of using our rescaling approach with $32$-laser or $16$-laser LiDAR rather than using zero-shot metric depth estimation. Thus, we observe on average $46$% improvement on $\delta_{1}$, $0.5$% on AbsRel and $47$% on RMSE relative to the other zero-shot methods. In contrast, the methods that have been fine-tuned on the same domain as the test set (see lines with (FT) in Tab. I and Tab. II) compare favorably to ours. These approaches benefit from an additional in-domain training which is often costly to perform due to dataset creation and training computation. We notice that our method is sensitive to the domain with on average $28$% and $13$% enhancements on indoor (including NYU V2) and outdoor (including KITTI) datasets, respectively, relative to other zero-shot methods whatever the metric. Regarding performance with 2D LiDAR, the same sensitivity to the domain is equally apparent. However, the results are a bit lower than zero-shot and fine-tuned metric depth estimation methods and lower than our other approaches, especially on outdoor datasets while remaining competitive on indoor datasets. Interestingly, we observe that increasing the number of beams from 16 to 32 does not necessarily improve the performance.
Rescaling with stereo
We show in our experiments on KITTI dataset that using stereo depth estimation with SGM [39] in order to rescale Depth Anything affine-invariant disparity predictions is more efficient than directly using SGM estimates. This improvement can be explained by the limited range of stereo methods that is exceeded by the depth of field in KITTI images. We carry out the same experiments on the DIML outdoor dataset [38] where the depth of field is much lower than in KITTI. Indeed, the results in Tab. V highlight that directly estimating depth with SGM [39] is better than using these estimates as a reference for rescaling even if DIML is part of the Depth Anything training set. Relative to the other methods we compare on the KITTI benchmark Tab. II, rescaling with stereo does not work as well as fine-tuned methods or rescaling with LiDAR 16 and 32 beams. However, rescaling with stereo performs on par with zero-shot metric depth estimation and is better than rescaling with SFM and LiDAR 1 beam.
Rescaling with structure-from-motion
Our experiments highlight that structure-from-motion techniques can provide reliable reference points for rescaling Depth Anything affine-invariant disparity predictions. More specifically, we show that results with SIFT [41] are slightly lower than the ones of other zero-shot monocular metric depth estimation methods, while using OmniGlue [42] increases performance relative to other zero-shot methods of the $\delta_{1}$, AbsRel and RMSE metrics on average by $7$%, $4$% and $19$%, respectively. Compared to the other approaches, rescaling with SFM performs better than rescaling with LiDAR 1 beam but worse than other methods that may however be more costly. This includes rescaling with LiDAR 16 and 32 beams or fine-tuned approaches that requires an extra training on a dataset that, in real cases, must be created.
Inference cost study
To complete our study, we analyze the inference cost of rescaling Depth Anything V1 affine-invariant disparity maps with other fine-tuned or zero-shot metric depth estimation methods in Tab. VI. For this purpose, we compare two informative and easy-to-access variables which are the number of parameters and the average inference runtime of these approaches for a single image. We notice that zero-shot methods are more costly at inference than our rescaling approach because they introduce extra modules such as a ConvGRU block in [31] or a camera module in [32] that returns a dense representation of the camera calibration. In contrast, fine-tuned methods have a reduced inference cost relative to ours thanks to a more complex training. Finally, as each image is processed independently, the computation overhead of our method can be even reduced when applied to a video stream. Indeed, since the neural network runs on the GPU and the RANSAC on the CPU, the inference of the neural network at step $t+1$ can overlap the execution of RANSAC at step $t$.
IV-D Comparison with Depth Anything V2
TABLE VII: Comparative study between Depth Anything V1 and Depth Anything V2 on the KITTI dataset [37].
Rescaling with Depth Anything KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $R^{2}$uparrow
LiDAR 1 V1 0.891 0.131 3.096 0.834
V2 0.733 0.193 7.053 0.721
LiDAR 16 V1 0.967 0.060 2.695 0.966
V2 0.961 0.067 2.829 0.954
LiDAR 32 V1 0.967 0.060 2.673 0.964
V2 0.962 0.067 2.815 0.953
We compare performance between Depth Anything V1 [2] and Depth Anything V2 [4] on the KITTI dataset in Tab. VII. We notice that Depth Anything V1 [2] slightly outperforms Depth Anything V2 [4] when used in our rescaling approach. Moreover, Depth Anything V1 has a higher coefficient of determination $R^{2}$ that corresponds to the proportion of the variance of the metric disparity that is explainable by the linear regression model parameterized with the $\alpha$ and $\beta$ that has been found. This means that while Depth Anything V2 manages to handle small details better than Depth Anything V1 (see Fig. 2), the latter predicts more accurate disparity maps within an affine transformation.
Figure 2: Qualitative comparison between Depth Anything V1 [2] and V2 [4]. From top to bottom: (1) image from the KITTI dataset [37], (2) the disparity map predicted by Depth Anything V1 and (2) the one of Depth Anything V2.
IV-E Ablation study
TABLE VIII: Ablation study to validate the need for estimating the rescaling parameters for each image
Rescaling with NYU V2
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
fixed parameters 0.787 0.142 1.240
Lidar 1 beam 0.939 0.063 0.652
Lidar 16 beams 0.976 0.039 0.454
[…]
KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
fixed parameters 0.908 0.106 3.976
Lidar 1 beam 0.813 0,152 6,032
Lidar 16 beams 0.966 0.060 2.755
[…]
SFM (SIFT) 0.910 0.103 3.686
SFM (OmniGlue) 0.921 0.103 3.179
[…]
TABLE IX: Ablation study to validate the use of the RANSAC algorithm.
Methods RANSAC KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
Ours rescaling w/ lidar 1 beam 0.897 0.113 4.871
0.891 0.131 3.096
Ours rescaling w/ lidar 16 beams 0.966 0.063 2.849
0.967 0.06 2.695
Ours rescaling w/ lidar 32 beams 0.966 0.063 2.852
0.967 0.06 2.673
Ours rescaling w/ SFM (SIFT) 0.001 0.926 19.053
0.893 0.103 3.92
Ours rescaling w/ SFM (OmniGlue) 0.859 0.107 4.012
0.925 0.093 3.562
Ours rescaling w/ stereo (SGM) 0.904 0.086 4.336
0.941 0.072 3.526
Tab. VIII compares dynamic rescaling i.e. an estimation of the parameters $\alpha$ and $\beta$ for each image to a static rescaling with unique parameters for all the images of a dataset. In the latter case, the parameters are the mean $\alpha$ and $\beta$ obtained with the best rescaling (here with LiDAR 16 beams). The results show that performing a rescaling for each image tends to provide better performance even if the camera does not change or if the image domain remains similar.
The ablation study in Tab. IX aims to justify the use of a RANSAC algorithm [6] when estimating the parameters of the affine transformation to recover the metric depth. We observe that the RANSAC is beneficial most of the time, especially when the 3D reference points are not provided by a LiDAR. Indeed, structure-from-motion or stereo-depth techniques are more likely to generate outliers that would disturb a vanilla linear regression but would be filtered out by the RANSAC.
V Conclusion
In this paper, we provide a low-cost method to estimate metric depth based on Depth Anything V1 [2] affine-invariant disparity predictions that are rescaled using 3D reference points provided by an additional sensor. Thus, the solution we propose is adaptable to any camera calibration and cheaper than traditional scale recovery approaches since it does not require any fine-tuning of the monocular depth estimation neural network. Moreover, by using predictions of Depth Anything V1 whose weights are publically available, we ensure generalization capacities to a large variety of image domains. Finally, our method is robust to noisy reference points or disparity maps, allowing the use of low-quality sensors for metric depth estimation at the lowest cost. To corroborate our claims we carry out experiments on standard depth estimation benchmarks that show that our approach is competitive with other zero-shot monocular depth estimation methods. Our future works will focus on confirming the interest of our approach by comparing it with a high-resolution LiDAR in the context of off-road navigation with a real robot.